OpenStudy (ujjwal):

The minimum number of vectors in different planes which can give zero resultant is 4.. But how??

OpenStudy (vincent-lyon.fr):

Take any two non-colinear vectors A and B: they define a plane. If any third vector C could make the sum 0, it would go A+B+C=0 or C=-A-B. C would then be a linear combination of A and B, which is a proof that C belongs to the plane defined by A and B. So you need a fourth vector, which is D = -A-B-C and is neither combination of (A,B), (A,C) or (B,C).

OpenStudy (ujjwal):

the vector D which you mention here would still lie in the same plane containing A,B and C.. Since, D=-A-B-C which is again a linear combination.. How do we get different planes here?

OpenStudy (anonymous):

A singe vector with a magnitude of zero produces a zero resultant. -- Two vectors with equal magnitudes and opposite directions produce a zero resultant. you'll need at least three. Think of them as being connected. To have a zero resultant, putting the vectors together head to tail should form a closed shape. The first vector can be in any direction. The second vector starts where the first ended, and extends in a different plane. The last vector starts from where the second ended and extends to the beginning of the first vector. The three end up making a triangle, which gives you a zero resultant

OpenStudy (ujjwal):

the three vectors will be in the same plane.. I am concerned about different planes..

OpenStudy (anonymous):

http://www.wiziq.com/tutorial/73966-Vector-Notes

OpenStudy (anonymous):

it may help u @ujjwal

OpenStudy (experimentx):